A particle has 37.0 joules of kinetic energy and 12.0 joules of gravitational potential energy at one point during its fall from the tree to the ground.An instant before striking to the ground,how much mechanical energy (rounded to the nearest joules) does the particle have? ignore air resistance.


1
Expert's answer
2021-06-29T08:18:24-0400
"T=K+U=37+12=49.0\ J"